regular milk has 4% butterfat. how many litres of regular milk must be mixed with 40L of milk with 1% butterfat to have milk with 2% butterfat?
Does someone know how to turn this into an equation? I would really appreciate it. thank you

Answer by stanbon(75887)

regular milk has 4% butterfat. how many litres of regular milk must be mixed with 40L of milk with 1% butterfat to have milk with 2% butterfat?
---------------------------------
Equation:
fat + fat = fat
0.04x + 0.01(40) = 0.02(x+40)
Multiply thru by 100 to get:
4x + 40 = 2x + 80
2x = 40
x = 20L (amount of regular mild needed in the mixture)
============================================================
Cheers,
Stan H.